OpenStudy (anonymous):

Please help!! Write an expression to describe a rule for the sequance. then find the 100th term in the sequance 3,10,17,24,31,38,... (1point) A: n+7;707 B: 7n;700 C: 7n-4;696 D: 6n-4;596

OpenStudy (anonymous):

How fast?

OpenStudy (anonymous):

you can take your time

OpenStudy (amorfide):

 let n be the term number term 1, term 2, term 3 you know the sequence is adding 7 each time so i wrote the 7 times table and underneath is the sequence itself now look at the table n=1 to get from 1 to 7, we multiply by 7 to get from 2 to 14 we multiply by 7 etc so we multiply n by 7 so we have 7n now compare the 7 with the first term 3 compare 14 with the second term 10 find the difference the difference is 7-3=4 14-10=4 so you have 7n-4 so your expression is an=7n-4